OpenStudy (anonymous):

-x + 3 = y for, y=x+b

OpenStudy (anonymous):

x=(3-b)/2

OpenStudy (anonymous):

fpr y=mx+b *

OpenStudy (anonymous):

x=(3-b)/(m+1)